# Frostman random variables, entropy inequalities, and applications

### v4

@@ 4.4 More general polynomials
$\begin{split}&d_{1}^{\prime\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{21}X+\xi_{22}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (80)
for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\Xi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\Xi)\geq 0$.
Now the theorem follows by combining (76) and (78)-(80), with $\epsilon_{4}=\epsilon_{2}/\max\{d_{1}+2d_{2},d_{1}^{\prime}+2d_{2}^{\prime},d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}\}$. ∎
For Theorem 1.4, the situation that $\phi(x,y)$ is diagonal has been proved by Lemma 4.8. The general case will be proved in the following form.
Theorem 1.4’ Suppose that $\phi(x,y)$ has the form (1). Let $0<s<1$, $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that the following statement holds: Let $X$ and $Y$ be two independent real random variables with ranges contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$
for all $n\in\mathbb{N}$.
@@
By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Theorem 1.3” provide us with some positive constant $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
@@
In the conclusions of Lemma 4.8 and Theorems 1.2’-1.4’, the implied constants in the terms $O_{d,s}(\log_{+}C)$, $O_{d,s}(\log_{+}C_{1}C_{2})$ or $O_{d,s}(\log_{+}c)$ can be independent of $s$ if we are only considering the case $s\gg 1$. We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
5 Proof of Theorem 1.4 with $s\leq 1/2$
The situation is more subtle for this case. It is hard to expect Theorem 2.38 holds true when the dimension gets smaller. We will use the upcoming theorem as a substitution. The core of the proof will follow the arguments from [24] by Shmerkin and Wang. However, there are enough differences that we cannot merely apply a single theorem of theirs as a black box, so to address this case, we will introduce various notions and results from their paper, and explicitly mention where our arguments diverge.
@@
Let $X$, $Y$ be independent random variables with ranges contained in [c,c][-c,c], with $c>0$. Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $0<s\leq 1/2$ and $C_{1},C_{2}\geq 1$. Let $(X^{\prime},Y^{\prime})$ be i.i.d. with $(X,Y)$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, we have
$H_{n}\big(|(X,Y)-(X^{\prime},Y^{\prime})|\big)\geq nt+O_{s}(\log n)+O_{s,t}(\log C_{1}C_{2})+O_{s,c}(1)$
@@
Here
$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$
@@
The rest of the proof of Theorem 1.4 for $s\leq 1/2$ will follow the same lines as in the case $s>1/2$. So we omit the details. To prove Theorem 5.1, we need the following variant for measures. For two measures $\mu^{\prime},\mu^{\prime\prime}$ whose supports are bounded regions on ${\mathbb{R}}^{2}$, we denote by $\Delta(\mu^{\prime},\mu^{\prime\prime})$ the distance measure between $\mu^{\prime}$ and $\mu^{\prime\prime}$, i.e.,
@@
for all sufficiently large $n$.
We emphasize that the logarithmic dependence on $C$ in the error term is essential. Because in the proof of the main theorem, we are working with infinitely many random variables satisfying (70), where the Frostman constants can be large compared to $n$, and are not uniform in $(x,y)$. Only an error of the form $\log C$ can be turned into entropy-terms as in (71). The main difficulty in this section, which prevents us from applying the results in [24] directly, is due to this reason.
Rather than repeating their arguments wholesale with minor modifications, we provide a detailed description of how to prove variants of their results tailored to our purposes, and combine these with a few calculations to prove Theorem 5.2.
@@
For simplicity, we denote $Z:=|(X,Y)-(X^{\prime},Y^{\prime})|$, and take $C=C_{1}C_{2}$. The random variables $\frac{X}{2c}+\frac{1}{2}$ and $\frac{Y}{2c}+\frac{1}{2}$ are both $\big(s,O((2c)^{s}C)\big)$-Frostman, with ranges contained in $[0,1]$. Moreover,
$H_{n}(|(X,Y)-(X^{\prime},Y^{\prime})|)=H_{n}\left(\left|\left(\frac{X}{2c}+\frac{1}{2},\frac{Y}{2c}+\frac{1}{2}\right)-\left(\frac{X^{\prime}}{2c}+\frac{1}{2},\frac{Y^{\prime}}{2c}+\frac{1}{2}\right)\right|\right)+O_{c}(1).$
Without loss of generality, we assume at the beginning of the proof that both $X$ and $Y$ are $(s,C)$-Frostman, with range contained in $[0,1]$. Our aim is to prove that
$H_{n}\big(Z\big)\geq nt+O_{s}(\log n)+O_{s,t}(\log C).$
Let $\gamma=\gamma(s,t)\in(0,1/3)$ be a parameter to be determined later. We apply Lemma 5.3 on the distribution measure of $X$, denoted by $\mu$. Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg_{s,t}C^{-1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2$.
@@
$H_{n}(Z)\geq\sum\limits_{i,j=1}^{3}P(E_{ij})H_{n}\big(Z_{E_{ij}}\big)=\sum\limits_{i,j=1}^{3}\mu(U_{i})\mu(U_{j})H_{n}\big(Z_{E_{ij}}\big).$
Since $X,Y,X^{\prime},Y^{\prime}$ are independent, the random variables $X_{U_{i}},Y,X_{U_{j}}^{\prime},Y^{\prime}$ are also independent, and $Z_{E_{ij}}\equiv|(X_{U_{i}},Y)-(X_{U_{j}}^{\prime},Y^{\prime})|$.
@@
$\geq 2abn(\varphi(2s)-\zeta)+(a^{2}+b^{2})ns+O_{s}(\log n)+O_{s,t}(\log C)$
$\geq 2(1/2-\gamma)^{2}(s+\varphi(2s)-\zeta)n+O_{s}(\log n)+O_{s,t}(\log C).$

### v5

@@ 4.4 More general polynomials
$\begin{split}&d_{1}^{\prime\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{21}X+\xi_{22}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}-1)\big(ns-\frac{1}{2}\log_{+}C_{1}C_{2}\big)+O_{\phi}(1)\end{split}$ (86)
for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\Xi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\Xi)\geq 0$. Now the theorem follows by combining (82) and (84)-(86), with
$\epsilon_{3}(d,\Xi,s)=\frac{\epsilon_{2}(d,s)}{\max\{d_{1}(\Xi)+2d_{2}(\Xi),\,d_{1}^{\prime}(\Xi)+2d_{2}^{\prime}(\Xi),\,d_{1}^{\prime\prime}(\Xi)+2d_{2}^{\prime\prime}(\Xi)\}}.$
Theorem 1.3”. Suppose that $\phi(x,y)$ has the form (2). Let $0<s<1$, $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{3}>0$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges both contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$
for all $n\in\mathbb{N}$. Here $\epsilon_{3}=\epsilon_{3}(d,s)$ when $\phi(x,y)$ is diagonal, and $\epsilon_{3}=\epsilon_{3}(d,\Xi,s)$ when $\phi(x,y)$ is non-diagonal.
@@
By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. So the conclusion of Theorem 1.3’ holds. , i.e.,
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{3})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
@@
In the conclusions of Lemma 4.8 and Theorems 1.2’, 1.3’, 1.3”, the implied constants in the terms $O_{d,s}(\log_{+}C)$, $O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})$ or $O_{d,s}(\log_{+}c)$ can be independent of $s$ if we are only considering the case $s\gg 1$. We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
To prove Theorem 1.3, we need a further property regarding the discretized joint entropy $H_{n}(X,Y)$.
Proposition 4.10.
Let $\phi(x,y)$ be given as in (2). Let $(X,Y)$ be a conditionally $(s,C;s,C)$-Frostman bivariate real random variable with range contained in $c,c]^{2}$, where $s,C,c>0$. Then
$H_{n}(X,Y)=H_{n}\big(X+Y,\,\phi(X,Y)\big)+O_{\phi,s,C,c}(1)$
for any $n\in\mathbb{N}$.
Proposition 4.10 will be proved within several lemmas, in which same notations and conditions are used. Let $F:\,{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ be defined by $F(x,y)=(x+y,\phi(x,y))$. Write $\mathfrak{J}(x,y)$ for its Jacobian matrix, and $J(x,y)$ for its Jacobian determinant. A direct computation yields
$J(x,y)$ $=\frac{\partial\phi}{\partial y}(x,y)-\frac{\partial\phi}{\partial x}(x,y)$
$=(\xi_{12}-\xi_{11})\rho_{1}^{\prime}(\xi_{11}x+\xi_{12}y)+(\xi_{22}-\xi_{21})\rho_{2}^{\prime}(\xi_{21}x+\xi_{22}y),$
which is a polynomial of degree at most d1d-1.
Lemma 4.11.
The Jacobian determinant $J(x,y)$ is not the zero polynomial.
Proof.
Suppose on the contrary that $J\equiv 0$. Consider the linear change of variables
$(u,v):=(\xi_{11}x+\xi_{12}y,\ \xi_{21}x+\xi_{22}y).$
Since $\Xi$ is invertible, $(x,y)\mapsto(u,v)$ is an invertible linear map, hence polynomial identities in $(x,y)$ are equivalent to polynomial identities in $(u,v)$. In $(u,v)$-coordinates, the equality $J\equiv 0$ appears as
$(\xi_{12}-\xi_{11})\,\rho_{1}^{\prime}(u)+(\xi_{22}-\xi_{21})\,\rho_{2}^{\prime}(v)\equiv 0.$
Such a sum can vanish identically only if both summands are constant and opposite. In view of $d\geq 2$, one sees that $\rho_{1}^{\prime}$ and $\rho_{2}^{\prime}$ are non-constant. It follows that $\xi_{12}=\xi_{11}$ and $\xi_{22}=\xi_{21}$. This contradicts the fact that $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$. ∎
By Lemma 4.11, the critical set
$\mathcal{Z}:=\{(x,y)\in{\mathbb{R}}^{2}:\ J(x,y)=0\}.$
is a (possibly reducible) algebraic curve. For any $\delta>0$, we denote by $\mathcal{N}_{\delta}(\mathcal{Z})$ the $\delta$-neighbourhood of the set $\mathcal{Z}$ in ${\mathbb{R}}^{2}$.
Lemma 4.12.
It satisfies that
$P\big((X,Y)\in\mathcal{N}_{\delta}(\mathcal{Z})\big)\ll_{\phi,c}C\delta^{s}.$
for all $\delta\in(0,1]$.
Proof.
Write $\mathcal{R}:=[-c,c]^{2}$ and $\Gamma:=\mathcal{Z}\cap\mathcal{R}$. Since $\Gamma$ is a compact semialgebraic set of dimension at most $1$, it admits a finite $C^{1}$-stratification. In particular, there exist integers $l=l(\phi),l^{\prime}=l^{\prime}(\phi)\geq 0$, $C^{1}$ arcs $\gamma_{1},\dots,\gamma_{l}\subseteq\mathcal{R}$, and points $p_{1},\dots,p_{l^{\prime}}\in\mathcal{R}$ such that
$\Gamma=\Big(\bigcup_{i=1}^{l}\gamma_{i}\Big)\cup\{p_{1},\dots,p_{l^{\prime}}\},$
and for each $i$ the arc $\gamma_{i}$ is the graph of a $C^{1}$ function either of the form $\{(x,f_{i}(x)):x\in I_{i}\}$ or $\{(g_{i}(y),y):y\in J_{i}\}$ on some compact intervals $I_{i},J_{i}$.
Let
$L:=1+\max_{1\leq i\leq l}\{\|f_{i}^{\prime}\|_{\infty}+\|g_{i}^{\prime}\|_{\infty}\}\ll_{\phi,c}1.$
Then for each graph arc we have the inclusions
$\mathcal{N}_{\delta}\big(\{(x,f_{i}(x))\}\big)\cap\mathcal{R}\subseteq\{(x,y)\in\mathcal{R}:\ |y-f_{i}(x)|\leq L\delta\},$
$\mathcal{N}_{\delta}\big(\{(g_{i}(y),y)\}\big)\cap\mathcal{R}\subseteq\{(x,y)\in\mathcal{R}:\ |x-g_{i}(y)|\leq L\delta\}.$
For an arc of the form $\{(x,f_{i}(x))\}$, it follows by the law of total expectation and the conditionally Frostman condition that
$P\big((X,Y)\in\mathcal{N}_{\delta}(\gamma_{i})\big)\leq P\big(|Y-f_{i}(X)|\leq L\delta\big)=\mathbb{E}\Big(P\big(|Y-f_{i}(X)|\leq L\delta\mid X\big)\Big)\leq C(L\delta)^{s}.$
Similarly, for an arc of the form $\{(g_{i}(y),y)\}$ the same bound holds. Finally, for each point $p_{j}=(p_{j,1},p_{j,2})$, we have
$P\big((X,Y)\in B(p_{j},\delta)\big)\leq P\big(|Y-p_{j,2}|\leq\delta\big)\leq C\delta^{s},$
since $(X,Y)$ is also marginally $(s,C;s,C)$-Frostman by Lemma 2.30.
Summing over the finitely many arcs and points yields
$P\big((X,Y)\in\mathcal{N}_{\delta}(\mathcal{Z})\big)\leq l\cdot C(L\delta)^{s}+l^{\prime}\cdot C\delta^{s}\ll_{\phi,c}C\delta^{s}.$
When $\mathcal{Z}$ is empty, it is not hard to prove Proposition 4.10. In the following, we assume $\mathcal{Z}$ to be non-empty, and apply the Łojasiewicz inequality (see [21, Theorem 2.12] for example) on our Jacobian determinant $J(x,y)$ and the distance function to $\mathcal{Z}$.
Lemma 4.13 (Łojasiewicz).
There exist constants $\gamma=\gamma(\phi,c)>0$ and $L=L(\phi,c)>0$ such that
$\text{dist}\big((x,y),\mathcal{Z}\big)\leq L|J(x,y)|^{1/\gamma}$
for all $(x,y)\in[-c,c]^{2}$.
Corollary 4.14.
It satisfies that
$P\big(|J(X,Y)|\leq 2^{-k}\big)\ll_{\phi,c}C2^{-ks/\gamma}$
for all $k\in\mathbb{N}$. In particular, if we define the integer-valued random variable $K:=\max\Big\{0,\ \big\lfloor-\log|J(X,Y)|\big\rfloor\Big\}$, then
$P(K\geq k)\ll_{\phi,c}C2^{-ks/\gamma}$
for all $k\geq 0$.
Proof.
By Lemma 4.13 and then Lemma 4.12, we have
$P(|J(X,Y)|\leq 2^{-k})\leq P\big((X,Y)\in\mathcal{N}_{L2^{-k/\gamma}}(\mathcal{Z})\big)\ll_{\phi,c}C(L2^{-k/\gamma})^{s}.$
The bound for $K$ follows since $\{K\geq k\}\subseteq\{|J(X,Y)|\leq 2^{-k}\}$. ∎
Next, let us consider the Jacobian matrix $\mathfrak{J}(x,y)$. In the region $(x,y)\in[-c-1,c+1]$, the operator norm satisfies $\|\mathfrak{J}(x,y)\|\leq\mathfrak{L}$ for some constant $\mathfrak{L}=\mathfrak{L}(\phi,c)>0$. Similarly, the Jacobian determinant $J$ is also $\mathfrak{L}^{\prime}$-Lipschitz function on $c-1,c+1]^{2}$ for some constant $\mathfrak{L}^{\prime}=\mathfrak{L}^{\prime}(\phi,c)$, i.e.,
$|J(x,y)-J(x^{\prime},y^{\prime})|\leq\mathfrak{L}^{\prime}\,|(x,y)-(x^{\prime},y^{\prime})|$ (87)
for all $(x,y),(x^{\prime},y^{\prime})\in[-c-1,c+1]^{2}$.
For fixed $(u,v)\in{\mathbb{R}}^{2}$, solving $F(x,y)=(u,v)$ means
$x+y=u,\qquad\phi(x,u-x)=v.$
The second equation is a univariate polynomial equation in $x$ of degree at most $d$. Hence for each fixed $u$ and each $v$, there are at most $d$ solutions for $x$ unless the polynomial in $x$ degenerates to an identically constant polynomial, which can occur only for finitely many $u$ and thus on a target set of Lebesgue measure 0. Therefore the fiber multiplicity function $N(u,v):=\#\{(x,y)\in[-c,c]^{2}:F(x,y)=(u,v)\}$ satisfies
$N(u,v)\leq d\quad\text{for Lebesgue-a.e.\ }(u,v)\in{\mathbb{R}}^{2}.$ (88)
Let $\mathcal{Q}_{n}$ denote the family of dyadic squares in ${\mathbb{R}}^{2}$ of side length $2^{-n}$. For $(u,v)\in{\mathbb{R}}^{2}$, let $Q_{n}(u,v)\in\mathcal{Q}_{n}$ be the unique dyadic square containing $(u,v)$.
Lemma 4.15.
For given $n\in\mathbb{N}$, let $B\in\mathcal{Q}_{n}$ be a dyadic square in the target $(u,v)$-plane. For $k\geq 0$, define
$E_{n,k}(B):=\Big\{(x,y)\in[-c,c]^{2}:\ F(x,y)\in B\ \text{and}\ 2^{-k-1}<|J(x,y)|\leq 2^{-k}\Big\}.$
Assume $k\leq n-\big\lceil\log\mathfrak{L}^{\prime}\big\rceil-5$. Then the number of dyadic input squares $Q\in\mathcal{Q}_{n}$ in the $(x,y)$-plane that intersect $E_{n,k}(B)$ has the upper bound
$\{Q\in\mathcal{Q}_{n}:Q\cap E_{n,k}(B)\neq\emptyset\}\ll_{\phi,c}2^{k}.$
Here the implied constant is independent of $n,k,B$.
Proof.
Let $\mathcal{S}$ be the set of dyadic input squares $Q\in\mathcal{Q}_{n}$ intersecting $E_{n,k}(B)$, and set $U:=\bigcup_{Q\in\mathcal{S}}Q$. Then the squares in $\mathcal{S}$ are disjoint and the area satisfies
$\mathcal{L}^{2}(U)=2^{-2n}(\#\mathcal{S}).$ (89)
Fix any $Q\in\mathcal{S}$. By definition, there exists $(x_{Q},y_{Q})\in Q\cap[-c,c]^{2}$ with $F(x_{Q},y_{Q})\in B$ and $|J(x_{Q},y_{Q})|>2^{-k-1}$. By (87) and the fact $k\leq n-\big\lceil\log\mathfrak{L}^{\prime}\big\rceil-5$, we have
$|J(x,y)-J(x_{Q},y_{Q})|\leq\mathfrak{L}^{\prime}\,|(x,y)-(x_{Q},y_{Q})|\leq\mathfrak{L}^{\prime}\,\sqrt{2}\,2^{-n}\leq 2^{-k-3}$
for all $(x,y)\in Q$, and then
$|J(x,y)|\geq|J(x_{Q},y_{Q})|-2^{-k-3}\geq 2^{-k-1}-2^{-k-3}\geq 2^{-k-2}.$
Hence $|J(x,y)|\geq 2^{-k-2}$ for all $(x,y)\in U$. Combining (89), one then deduce that
$\iint_{U}|J(x,y)|\,dxdy\geq 2^{-k-2}\mathcal{L}^{2}(U)\gg 2^{-2n-k}(\#\mathcal{S}).$ (90)
Next, since $\|\mathfrak{J}\|\leq\mathfrak{L}$ on $c-1,c+1]^{2}$, the image of each $Q$ has diameter at most $\mathfrak{L}\,\sqrt{2}2^{-n}$. Because $F(x_{Q},y_{Q})\in B$, it follows that $F(Q)$ is contained in a fixed enlargement $B^{+}$ of $B$ by an $O(\mathfrak{L}2^{-n})$ margin, hence $F(U)\subseteq B^{+}$, with the area satisfying
$\mathcal{L}^{2}(B^{+})\ll_{\phi,c}2^{-2n}.$ (91)
Now applying the area formula to the Lipschitz map $F$ as a change of variable on the measurable set $U$, combining (88) and (91), we get
$\iint_{U}|J(x,y)|\,dxdy\leq\iint_{F(U)}N(u,v)\,dudv\leq d\,\mathcal{L}^{2}(F(U))\leq d\,\mathcal{L}^{2}(B^{+})\ll_{\phi,c}2^{-2n}.$ (92)
Putting (90) and (92) together, one concludes $S\ll_{\phi,c}2^{k}$. The lemma then follows. ∎
Proof of Proposition 4.10.
For simplicity, we write
$Z_{n}:=D_{n}(X,Y),\qquad W_{n}:=D_{n}\big(X+Y,\phi(X,Y)\big).$
By mean value inequality, one can see that $W_{n}$ is an $\mathfrak{L}$-Lipschitz function of $Z_{n}$, where, as previous, $\mathfrak{L}=\mathfrak{L}(\phi,c)$ is an upper bound for the operator norm of the Jacobian matrix $\mathfrak{J}(x,y)$ in the region $c-1,c+1]^{2}$. Lemma 2.16 shows that
$H_{n}(W_{n})\leq H_{n}(Z_{n})+O_{\phi,c}(1).$
For the opposite direction, in view of (6), it is sufficient to prove that
$H(Z_{n}|W_{n})=O_{\phi,s,C,c}(1).$ (93)
Recall the random variable $K$ from Corollary 4.14, we will use the inequality
$H(Z_{n}\mid W_{n})\leq H(K)+H(Z_{n}\mid W_{n},K)$ (94)
First, by Corollary 4.14 one has
$p_{k}:=P(K=k)\leq P(K\geq k)\leq M2^{-ks/\gamma},\quad(k\geq 0)$
for some constant $M=M(\phi,C,c)\geq 1$. Here $\gamma$ is the constant in Lemma 4.13 that only depends on $\phi$ and $c$. Now calculation gives
$\mathbb{E}(K)=\sum_{k\geq 1}P(K\geq k)\leq M\sum_{k\geq 1}2^{-ks/\gamma}\ll_{\phi,s,C,c}1.$ (95)
Moreover, take $k_{1}=\lceil(\gamma/s)\log(8M)\rceil$. For any $k\geq k_{1}$, one has $p_{k}<1/e$, so by monotonicity one has
$p_{k}\log p_{k}\leq M2^{-ks/\gamma}(-\log M+ks/\gamma)\leq(Mks/\gamma)\,2^{-ks/\gamma}.$
It follows that
$\sum_{k\geq k_{1}}p_{k}\log p_{k}\leq(Ms/\gamma)\sum\limits_{k\geq k_{1}}k2^{-ks/\gamma}\ll_{\phi,s,C,c}1.$
Moreover,
$\sum_{0\leq k<k_{1}}p_{k}\log p_{k}\leq k_{1}\ll_{\phi,s,C,c}1.$
since $p_{k}\log p_{k}\leq 1$. It follows that
$H(K)=-\sum\limits_{k\geq 0}p_{k}\log p_{k}\ll_{\phi,s,C,c}1.$ (96)
Second, we fix $n$ and condition on the event $\{W_{n}=(u,v),\,K=k\}$. The value $W_{n}=(u,v)$ fixes the target dyadic square $B:=Q_{n}(u,v)$ intersecting the range of $(X+Y,\phi(X,Y))$.
When $k\leq n-\big\lceil\log\mathfrak{L}^{\prime}\big\rceil-5$, Lemma 4.15 shows that the number of dyadic input squares $Q\in\mathcal{Q}_{n}$, which can intersect the range of $(X,Y)$ while still being consistent with $F(X,Y)\in B$ and the Jacobian bound, is at most $O_{\phi,c}(2^{k})$. Therefore,
$H(Z_{n}\mid W_{n}=w,K=k)\leq\log\big(O_{\phi,c}(2^{k})\big)=k+O_{\phi,c}(1).$
When $k>n-\big\lceil\log\mathfrak{L}^{\prime}\big\rceil-5$, we use the trivial bound for the discretized variables $Z_{n}$ in $c,c]^{2}$:
$H(Z_{n}\mid W_{n}=w,K=k)\leq H(Z_{n})\leq 2n+O_{c}(1)\leq 2k+O_{\phi,c}(1),$
where we have used Lemma 2.4(iv), (8) and the fact that $n<k+\big\lceil\log\mathfrak{L}^{\prime}\big\rceil+5$. Combining the two regimes, we see that uniformly for all $k\geq 0$,
$H(Z_{n}\mid W_{n}=w,K=k)\leq 2k+O_{\phi,C,c}(1).$
Taking the expectation over $K$, we obtain by (95) that
$H(Z_{n}\mid W_{n},K)\leq 2\mathbb{E}(K)+O_{\phi,C,c}(1)=O_{\phi,s,C,c}(1).$ (97)
Combing (94), (96) and (97), we obtain (93). The proposition then follows. ∎
Finally, we complete the proof of Theorem 1.3.
Proof of Theorem 1.3.
Since $(X,Y)$ is conditionally $(s,C;s,C)$-Frostman, an application of Proposition 4.10 leads to
$\max\{H_{n}(X+Y),\,\phi(X,Y)\}\geq\frac{1}{2}H(X,Y)+O_{\phi,s,C,c}(1).$
Inserting the above inequality to the first-term in the maxima in the conclusion of Theorem 1.3’ with the conditionally $(s,C;s,C)$-Frostman random variable $(X,Y)$, we obtain
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{3})+O_{\phi,s,C,c}(1).$
Theorem 1.3 then follows. ∎
5 Proof of Theorems 1.2 and 1.3 with $s\leq 1/2$
When $s\leq 1/2$, as noted in the introduction, the distance–entropy input used for $s>1/2$ cannot yield a gain beyond the Frostman exponent. In this range, we instead use a product-structure distance mechanism, following Shmerkin and Wang [35], and this is the reason for the independence hypothesis in Theorem 5.1.
Define the increasing function
$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$
@@
Let $X$, $Y$ be independent random variables with ranges contained in [c,c][-c,c], where $c>0$. Suppose that $(X,Y)$ is marginally $(s,C;s,C)$-Frostman, where $0<s\leq 1/2$ and $C>0$. Let $(X^{\prime},Y^{\prime})$ be i.i.d. with $(X,Y)$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, we have
$H_{n}\big(|(X,Y)-(X^{\prime},Y^{\prime})|\big)\geq nt+O_{s}(\log n+\log_{+}c)+O_{s,t}(\log_{+}C+1)$
@@
The proof has three parts. We first reduce Theorem 5.1 to an entropy lower bound for suitable distance measures, which is stated as Theorem 5.2. Then, we prove Theorem 5.2 by adapting the robust projection and thin-tube arguments of Shmerkin and Wang to our entropy setting, while keeping explicit logarithmic dependence on the Frostman constant. Finally, we return to the random-variable formulation and deduce Theorem 5.1.
@@
The rest of the proof of Theorems 1.2 and 1.3 for $s\leq 1/2$ will follow the same lines as in the case $s>1/2$. So we omit the details. To prove Theorem 5.1, we need the following variant for measures. For two measures $\mu^{\prime},\mu^{\prime\prime}$ whose supports are bounded regions on ${\mathbb{R}}^{2}$, we denote by $\Delta(\mu^{\prime},\mu^{\prime\prime})$ the distance measure between $\mu^{\prime}$ and $\mu^{\prime\prime}$, i.e.,
@@
for all $n\in\mathbb{N}$.
We emphasize that the logarithmic dependence on $C$ in the error term is essential. Because in the proof of the main theorem, we are working with infinitely many random variables satisfying (63) or (76), where the Frostman constants can be large compared to $n$, and are not uniform in $(z_{1},z_{2})$ or $(x,y)$. Only an error of the form $\log C$ can be turned into entropy-terms as in (68) or (77), respectively. The main difficulty in this section, which prevents us from applying the results in [35] directly, is due to this reason.
@@
If $c>1$, then we consider the independent random variables $\frac{X}{2c}+\frac{1}{2}$ and $\frac{Y}{2c}+\frac{1}{2}$. They are both $\big(s,O((2c)^{s}C)\big)$-Frostman, and have range contained in $[0,1]$. Also,
$H_{n}(|(X,Y)-(X^{\prime},Y^{\prime})|)\geq H_{n}\left(\left|\left(\frac{X}{2c}+\frac{1}{2},\,\frac{Y}{2c}+\frac{1}{2}\right)-\left(\frac{X^{\prime}}{2c}+\frac{1}{2},\,\frac{Y^{\prime}}{2c}+\frac{1}{2}\right)\right|\right)+O(1).$
Then we can work with $\big(\frac{X}{2c}+\frac{1}{2},\,\frac{Y}{2c}+\frac{1}{2}\big)$ instead of $(X,Y)$. Writing $Z:=|(X,Y)-(X^{\prime},Y^{\prime})|$, we assume without loss of generality at the beginning of the proof that the ranges of $X$ and $Y$ are both contained in $[0,1]$. Our aim is to prove that
$H_{n}\big(Z\big)\geq nt+O_{s}(\log n)+O_{s,t}(\log C+1).$
Let $\gamma=\gamma(s,t)\in(0,1/3)$ be a parameter to be determined later. We apply Lemma 5.3 on the distribution measure of $X$, denoted by $\mu$. Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg(\gamma/C)^{1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2$.
@@
$H_{n}(Z)\geq\sum\limits_{i,j=1}^{3}P(E_{ij})H_{n}\big(Z_{E_{ij}}\big)=\sum\limits_{i,j=1}^{3}\mu(U_{i})\mu(U_{j})H_{n}\big(Z_{E_{ij}}\big),$
where $Z_{E_{ij}}$ is the random variable such that $Z_{E_{ij}}\equiv(Z|E_{ij})$. Since $X,Y,X^{\prime},Y^{\prime}$ are independent, the random variables $X_{U_{i}},Y,X_{U_{j}}^{\prime},Y^{\prime}$ are also independent, and $Z_{E_{ij}}\equiv|(X_{U_{i}},Y)-(X_{U_{j}}^{\prime},Y^{\prime})|$. Here $X_{U_{i}}$ and $X_{U_{j}}$ denote the random variables such that $X_{U_{i}}\equiv(X|X\in U_{i})$ and $X^{\prime}_{U_{j}}\equiv(X^{\prime}|X^{\prime}\in U_{j})$.
@@
$\geq 2abn(\varphi(2s)-\zeta)+(a^{2}+b^{2})ns+O_{s}(\log n)+O_{s,t}(\log C+1)$
$\geq 2(1/2-\gamma)^{2}(s+\varphi(2s)-\zeta)n+O_{s}(\log n)+O_{s,t}(\log C+1).$